A grain storage warehouse has a total of 30 bins. Some hold 20 tons of grain each, and the rest hold 15 tons of grain each. How many of each type of bin are there if the capacity of the warehouse is 510?
Answer by mananth(16949)

A grain storage warehouse has a total of 30 bins. Some hold 20 tons of grain each, and the rest hold 15 tons of grain each. How many of each type of bin are there if the capacity of the warehouse is 510?
...
let number of 20 T bins be x
& 15 T bins be y
x+y = 30 ...............................1
20x+15y = 510 .........................2
...
multiply 1 by -20
-20x-20y=-600
add this to equation 2
-20x+20x+15y-20y=510-600
-5y=-90
/-5
-5y/-5 = -90/-5
y=18 ............. 15T bins
...
plug the value of y in 1
x+18=30
-18 -18
x=30-18
x= 12 , the number of 20 T bins.
...
m.ananth@hotmail.ca